OpenStudy (anonymous):

The table below represents the velocity of a car as a function of time: The average rate of change of the function between X= 15 to X= 29 is ____ m/s to the second power and represents the car's acceleration. Please help!

OpenStudy (danjs):

for the average rate of chainge in velocity, you can imagine a line through those 2 endpoints. The slope of that is the average change over that interval

OpenStudy (anonymous):

Oh so it's 2?

OpenStudy (danjs):

points (time, distance) .. (t , X)

OpenStudy (danjs):

not sure , how much time has elapsed from X=15 to X=29 on your table

OpenStudy (anonymous):

42

OpenStudy (danjs):

ok, then the slope is the change in X over the change in time (29 - 15) / 42

OpenStudy (anonymous):

So it's 3????

OpenStudy (anonymous):

@Danjs 3?

OpenStudy (danjs):

what are the two points X=15,t=? X=29, t=?

OpenStudy (anonymous):

I have no idea/:

OpenStudy (danjs):

just the slope, difference in the X, over difference in the time t (29-15) / 42 = 14/42, simploify to reduced fraction

OpenStudy (danjs):

i don't either, you have the table not me.. :)

OpenStudy (anonymous):

Ohhhh X= 15 f(X)= 45 X= 29 f(X)= 87

OpenStudy (danjs):

ok then x is the time here, and f(x) is the velocity average accel = change in velocity / change in time = (87 - 45) / (29 - 15) m/s^2

OpenStudy (anonymous):

So what should I put as the answer?

OpenStudy (danjs):

= 42 / 14 = 6 / 2 = 3 m/s^2

OpenStudy (anonymous):

So 3? Thank you so much!

OpenStudy (danjs):

welcome